OpenStudy (anonymous):

The scores listed below are the results from students taking a dexterity test. What is the 75th percentile of these scores? 1, 1, 1, 1.5, 2, 2, 2, 2.5, 3 A) 2 B) 2.25 C) 2.5 im torn between a and b

OpenStudy (anonymous):

Lp = (n +1) 75/100 Lp = 9 *75 / 100 Lp = 6.75th observation, using 7th observation. therefore, the answer is 2...

OpenStudy (anonymous):

u can see here http://answers.yahoo.com/question/index?qid=20110413205732AA5a97t

OpenStudy (anonymous):

@maariee did u understand

OpenStudy (anonymous):

yes thanks so much! :)